OpenStudy (anonymous):

how do i find the real solutions of [ cot theta=0 ]

OpenStudy (anonymous):

at 0? maybe im not quite sure

OpenStudy (anonymous):



OpenStudy (anonymous):

thats all it gives me

OpenStudy (mertsj):

cotangent is cosine/sine If the cosine is 0, the fraction is 0 so find the values for which the cosine is 0

OpenStudy (anonymous):

ok thank you. it would be pi/2 in the interval 0,2pi

OpenStudy (mertsj):

And 3pi/2

OpenStudy (anonymous):

ahhh yeah 3pi/2. thank you

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

what about cot^2x = 3

OpenStudy (anonymous):

i figure it to be cot x = radical 3 im not sure where to go from there

OpenStudy (mertsj):

\[\cot ^2x=3\] \[\cot x=\pm \sqrt{3}\] \[\tan x=\frac{1}{\cot x}=\pm \frac{1}{\sqrt{3}}=\pm \frac{\sqrt{3}}{3}\]

OpenStudy (mertsj):

So find the values of x that have that tangent

OpenStudy (anonymous):

30 or pi/6